A compound of only carbon and bromine consists of 10.15% carbon and has a molar mass of 707.70 g/mol. Determine the molecular formula of this compound.


Please answer with correct number of sig figs, scientific notation, and btw this question has two parts.



1
Expert's answer
2021-02-11T04:46:33-0500

Assuming that you have 100g of the compound; Mass of Carbon=10.15 and mass of Br=(100-10.150) =89.85

C Br

Mass 10.15 89.85


RAM 16 35


Moles "\dfrac{10.15}{12}" "\dfrac{89.85}{35}"


0.84583 2.56714


Mole ratios "\dfrac{0.84583}{0.84583}" "\dfrac{2.56714}{0.84583}"


1 : 3


Empirical formula = CBr3


(CBr3)n = 707.7


(12+105)n =707.7


"\dfrac{117n}{117}=\dfrac{707.7}{117}"


n=6.05

MF=(CBr3)6 = C6Br12